OpenStudy (anonymous):

y varies inversely as twice x. When x = 3, y = 8. Find y when x = 2.

OpenStudy (saifoo.khan):

y = k/2x

OpenStudy (saifoo.khan):

find the value of k.

OpenStudy (anonymous):

then solve for ur x

OpenStudy (saifoo.khan):

8*2(3) = k k = 48 now: y = 48/2x

OpenStudy (saifoo.khan):

y = 48/2(2) y = 48/4 y = 12.